a mixture of 75 cm3of oxygen and 12.5 cm3 of a gaseous hydrocarbon h were exploded in an eudiometer. after cooling to room temperature, 50 cm3 of gas was left and when this was shaken with koh solution, 12.5 cm3 of oxygen remained. given that all the volumes were measured at same temperature and pressure, calculate the molar mass of the hydrocarbon h.
1
Expert's answer
2020-06-11T10:26:48-0400

2CnH2n+2+3n+1 O2= 2n CO2+2n+2H2O

1.we find the amount of oxygen that reacts

75-12.5=62.5sm3

2.we can determine the proportion of hydrocarbons

2mol hydrocarbons 3n+1 oxygen

12.5sm3 hydrocarbons 62.5sm3 oxygen

we get the value of n from the proportion

3n+1=62.5*2/12.5

3n+1=10

3n=9

n=3

hence the formula of the hydrocarbon is C3H8

3.we calculate the molar mass of C3H8

12*3+1*8=44g/mol